# Manifold Learning via Foliations and Knowledge Transfer

### v1

@@ 5 Conclusions
We propose to complement the notion of learning manifolds with the more general one of learning foliations. The (integrable) distribution $\mathcal{D}$, defined via the data information matrix (DIM), a generalization of the Fisher information matrix to the space of data, indeed allows for the partition of the data space according to the leaves of such a foliation via the Frobenius theorem. Examples and experiments show a correlation of data points with the leaves of the foliation: moving according to the distribution $\mathcal{D}$, i.e. along a leaf, the model gives a meaningful label, while moving in the orthogonal directions leads to greater and greater classification errors. The learning foliation is however both singular (drop in rank) and non smooth. We prove that singular points are contained into a set of measure zero, hence making the learning foliation significant in the data space. We show that points in the dataset the model was trained with have lower DIM eigenvalues, so that the distribution $\mathcal{D}$ allows successfully to determine whether a sample of points belongs or not to the dataset used for training. We make such explicit comparison with similar datasets (i.e. MNIST versus FashionMNIST, KMNIST etc). Then, we use the lowest eigenvalue of the DIM to measure the distance between data sets. We test our proposed distance by retraining our model on datasets belonging to the same data spaces and checking the validation accuracy. Our results are not quantitatively conclusive in this regard, but show a great promise as a first step to go beyond the manifold hypothesis and exploiting the theory of singular foliations to perform dimensionality reduction and knowledge transfer.

### v2

@@ 5 Conclusions
We propose to complement the notion of data manifolds with the more general one of data foliations. The (integrable) distribution $\mathcal{D}$, defined via the data information matrix (DIM), a generalization of the Fisher information matrix to the space of data, indeed allows for the partition of the data space according to the leaves of such a foliation via the Frobenius theorem. Examples and experiments show a correlation of data points with the leaves of the foliation: moving according to the distribution $\mathcal{D}$, i.e. along a leaf, the model gives a meaningful label, while moving in the orthogonal directions leads to greater and greater classification errors. The data foliation is however both singular (drop in rank) and non smooth. We prove that singular points are contained into a set of measure zero, hence making the data foliation significant in the data space. We show that points in the dataset the model was trained with have lower DIM eigenvalues, so that the distribution $\mathcal{D}$ allows successfully to determine whether a sample of points belongs or not to the dataset used for training. We make such explicit comparison with similar datasets (i.e. MNIST versus FashionMNIST, KMNIST etc). Then, we use the average magnitude of DIM’s eigenvalues to understand the distance between data sets. We test our proposed distance by retraining our model on datasets belonging to the same data spaces and checking the validation accuracy. Our results are not quantitatively conclusive in this regard, but they represent a "proof of concept", to encourage further investigation. We believe that our empirical examples show a great promise as a first step to go beyond the manifold hypothesis and exploiting the theory of singular foliations to perform dimensionality reduction and knowledge transfer.
Acknowledgments
This research was supported by Gnsaga-Indam, by COST Action CaLISTA CA21109, HORIZON2020 CaLIGOLA MSCA-2021-SE-01-101086123 MSCA-DN CaLiForNIA-101119552, PNRR MNESYS, PNRR National Center for HPC, Big Data and Quantum Computing, PNRR SymQUSEC, INFN Sezione Bologna and PEPR “ANR-23-PEIA-0004”.
Appendix A Neural Network Architecture
In this section, you can find on table˜3 and figure˜7 the architecture of the ReLU network trained on MNIST used for the experiment in section˜4.
Table 3: Architecture of the neural network trained on MNIST.
No. Layers (sequential)
(0): Conv2d(1, 32, kernel_size=(3, 3), stride=(1, 1))
(1): ReLU()
(2): Conv2d(32, 64, kernel_size=(3, 3), stride=(1, 1))
(3): ReLU()
(4): AvgPool2d(kernel_size=2, stride=2, padding=0)
(5): Flatten()
(6): Linear(in_features=9216, out_features=128, bias=True)
(7): ReLU()
(8): Linear(in_features=128, out_features=10, bias=True)
(9): Softmax()
Figure 7: The structure of our CNN - picture created with [lenail].
Appendix B Proof of proposition˜3.3
Note that these computations can also be found in [gf2021, tronCartanMovingFrames2024] under different forms.
First, recall that
$\left[u^{k}\partial_{k},v^{l}\partial_{l}\right]=\left(u^{k}\frac{\partial v^{l}}{\partial x^{k}}-v^{k}\frac{\partial u^{l}}{\partial x^{k}}\right)\partial_{l}.$
Thus, if $1\leq a,b\leq c-1$, $u^{k}=\frac{\partial p_{a}}{\partial x^{k}}$ and $v^{l}=\frac{\partial p_{b}}{\partial x^{l}}$ we get:
$\left[u,v\right]$ $=\left(\sum_{k}\frac{\partial p_{a}}{\partial x^{k}}\frac{\partial^{2}p_{b}}{\partial x^{k}\partial x^{l}}-\sum_{k}\frac{\partial p_{b}}{\partial x^{k}}\frac{\partial^{2}p_{a}}{\partial x^{k}\partial x^{l}}\right)\partial_{l}$
$=H\left(p_{b}\right)u-H\left(p_{a}\right)v$ (16)
with $H(p_{a})$ the Hessian matrix of $p_{a}=N(x)_{a}$.
Let $s$ represent the score vector, i.e. the output of the neural network right before going through the softmax function. Then, for all $i,j=1,\ldots,c$,
$\partial_{j}p_{i}=\sum_{k}p_{i}\left(\delta_{ik}-p_{k}\right)\partial_{j}s_{k}.$ (17)
This is straightforward from the definition of the softmax function.
We can then express the Hessian matrix of $p_{a}$ with respect to $s$ and its derivative as follow.
$H\left(p_{a}\right)_{ij}$ $=\partial_{i}\left(\partial_{j}p_{a}\right)=\partial_{i}\left(\sum_{k}p_{a}\left(\delta_{ik}-p_{k}\right)\partial_{j}s_{k}\right)$
$=\partial_{i}p_{a}\sum_{k}\left(\delta_{ak}-p_{k}\right)\partial_{j}s_{k}+p_{a}\sum_{k}\partial_{i}\left(\left(\delta_{ak}-p_{k}\right)\partial_{j}s_{k}\right)$
$=\partial_{i}p_{a}\sum_{k}\left(\delta_{ak}-p_{k}\right)\partial_{j}s_{k}+p_{a}\sum_{k}\left(\delta_{ak}-p_{k}\right)H\left(s_{k}\right)-p_{a}\sum_{k}\partial_{i}p_{k}\partial_{j}s_{k}$
$=\sum_{k}\left(\left(\delta_{ak}-p_{k}\right)\partial_{i}p_{a}-p_{a}\partial_{i}p_{k}\right)\partial_{j}s_{k}+p_{a}\sum_{k}\left(\delta_{ak}-p_{k}\right)H\left(s_{k}\right).$
In the case of ReLU activation functions, the second order derivative of the score $s$ vanishes almost everywhere [^7] This question is the focus of theorem 3.6.. We can thus safely discard the last term.
Now, we compute $H(p_{a})\nabla_{x}p_{b}$ by plugging in the above expression, taking advantage of $H(p_{a})$’s symmetry:
$\left(H\left(p_{a}\right)\nabla_{x}p_{b}\right)_{j}$ $=\sum_{i}\left(\sum_{k}\left(\left(\delta_{ak}-p_{k}\right)\partial_{j}p_{a}-p_{a}\partial_{j}p_{k}\right)\partial_{i}s_{k}\right)\partial_{i}p_{b}$
[…]
$=\left(\sum_{k,i}\left(\delta_{ak}-p_{k}\right)\partial_{i}s_{k}\partial_{i}p_{b}\right)\partial_{j}p_{a}-p_{a}\sum_{k,i}\left(\partial_{i}s_{k}\partial_{i}p_{b}\right)\partial_{j}p_{k}.$
Therefore,
$H(p_{a})\nabla_{x}p_{b}=\left(\sum_{k,i}\left(\delta_{ak}-p_{k}\right)\partial_{i}s_{k}\partial_{i}p_{b}\right)\nabla_{x}p_{a}-p_{a}\sum_{k,i}\left(\partial_{i}s_{k}\partial_{i}p_{b}\right)\nabla p_{k}.$
This conclude the proof by showing that, for ReLU activation functions, we have $H(p_{a})\nabla_{x}p_{b}\in\mathrm{span}_{i=1,\ldots,c}\{\nabla_{x}\ln p(y_{i}|x,w)\}$ Thus the distribution $\mathcal{D}$ is locally involutive and Frobenius theorem applies.